How do solve step by step x^2=40? I am suppposed to solve the equation.
Thanks

Found 2 solutions by Alan3354, unlockmath:
Answer by Alan3354(69443)

There aren't many steps.
x^2 = 40
x = +sqrt(40) = +2sqrt(10)
x = -sqrt(40) = -2sqrt(10)
---------------------------
Use a calculator, --> x =~ 6.324555

Answer by unlockmath(1688)

Hello,
The whole key to algebra equations is get the variable "x" all alone. So when we have an equation like this: x^2=40
The only way to get it to be "x" is square root both sides. When we sq rt x^2 it just becomes "x". But we need to sq rt the other side too so to keep it balanced.
(sq rt)x^2=(sq rt)40
x=6.3246 Rounded off (Also a minus 6.32)
Now we can check this to see if it's right by putting the answer back into the original equation.
RJ
Check out a new book I wrote at:
www.math-unlock.com